OpenStudy (binary3i):

g(x)g(y)=g(x)+g(y)+g(xy)-2. if g(2)=5 then g(3)=?

OpenStudy (binary3i):

can any one

OpenStudy (mathmate):

Try finding g(1) first. Given g(2)=5 g(2)g(1)=g(2)+g(1)+g(2*1)-2 => 5g(1)=5+g(1)+5-2 => g(1)=4 Now g(1)g(3)=g(1)+g(3)+g(3*1)-2 => 4g(3)=4+g(3)+g(3)-2 => g(3)=1

OpenStudy (asnaseer):

@mathmate beat me to it :-)

OpenStudy (mathmate):

Try finding g(1) first. Given g(2)=5 g(2)g(1)=g(2)+g(1)+g(2*1)-2 => 5g(1)=5+g(1)+5-2 => g(1)=2 (mistake in previous post) Now g(1)g(3)=g(1)+g(3)+g(3*1)-2 => 2g(3)=2+g(3)+g(3)-2 => 0=0 Sorry, it doesn't work.

OpenStudy (mathmate):

asnaseer, it didn't work out. Hope you have another idea! :)

OpenStudy (asnaseer):

this is what I was about to post previously: if you let x=y=2 you will get:\[g(2)*g(2)=g(2)+g(2)+g(4)-2\]\[25=10+g(4)-2\]\[g(4)=17\]maybe you can use similar techniques to solve for g(3)?

OpenStudy (asnaseer):

we can also find g(0) as follows: g(0)*g(2)=g(0)+g(2)+g(0)-2 5g(0)=2g(0)+5-2=2g(0)+3 g(0)=1

OpenStudy (asnaseer):

this leads to: g(0) = 1 g(1) = 2 g(2) = 5 g(3) = x g(4) = 17 which could be interpreted as a sequence: 1, 2, 5, x, 17 difference between each term looks like the odd numbers so: 1+1 = 2 2+3 = 5 5+5 = 10 10+7 = 17 so I would say g(3)=10

OpenStudy (mathmate):

From g(2)*g(4), we find g(8)=65. Also, from g(4)=17, we get: 17+(9+11+13+15)=65! However, it's not a proof.